# Real-time Motion Planning for autonomous vehicles in dynamic environments

### v1

@@ Abstract
Index Terms:
Autonomous Vehicles, Dynamic obstacles, Obstacle avoidance, Global planning, Local planning, Timed elastic band, trajectory density
I Introduction
@@ II Methodology
II-A Proposed algorithm for global planning
@@ II-B $A^{*}$ planner
II-C Gradient descent method
@@
II-D Proposed algorithm for local planning
II-E Setting the trajectory density
@@
II-F Obstacle dynamics in local planning
@@
TABLE I: System Augmented State
@@
$x\hat{\ }_{\text{aug}}(n|n-1)=F_{\text{aug}}(n)x\hat{\ }_{\text{aug}}(n-1|n-1)\\ \mathrm{\Sigma}_{\text{aug}}(n|n-1)=F_{\text{aug}}(n)\mathrm{\Sigma}_{\text{aug}}(n-1|n-1)F_{\text{aug}}^{T}(n)\\ +G_{\text{aug}}(n)Q_{\text{aug}}(n)G_{\text{aug}}^{T}(n)\\ y\hat{(n)}=Z_{\text{aug}}(n)-H_{\text{aug}}(n)x\hat{\ }_{\text{aug}}(n|n-1)\\ K_{\text{aug}}(n)=[\mathrm{\Sigma}_{\text{aug}}(n|n-1)H_{\text{aug}}^{T}(n)\\ +G_{\text{aug}}(n)T_{\text{aug}}(n)]R_{\text{aug}}^{-1}(n)\\ x\hat{\ }_{\text{aug}}(n|n)=x\hat{\ }_{\text{aug}}(n|n-1)+K_{\text{aug}}(n)y\hat{(n)}\\ \mathrm{\Sigma}_{\text{aug}}(n|n)=\mathrm{\Sigma}_{\text{aug}}(n|n-1)\\ -K_{\text{aug}}(n)H_{\text{aug}}(n)\mathrm{\Sigma}_{\text{aug}}(n|n-1)$ (10)
@@
The desired dynamic system is modeled using position, speed and acceleration. Now it is assumed that the acceleration, which is larger than the order, is constant in specific time intervals that are independent from each other and changes at the end of each interval. In other words, the acceleration jumps in each time interval and is modeled as below.
$f(x)=Fx+\Gamma w$ (12)
where $\Gamma$ is the system noise gain and $w$ is the desired continuous piece acceleration. The transfer matrix of the system is also defined as follows:
@@
In a period of time, the velocity changes as much as $w(t)\Delta t$ and the position changes as much as $w(t)\left(\Delta t\right)^{2}/2$. As a result, it can be found that
@@
III Experiments
Fig.2 illustrates the scenario involving a parking lot, showcasing two planning methods: normal $A^{*}$ and gradient descent. Comparing the paths generated by these methods reveals that the gradient descent approach yields a notably smoother and optimal path compared to the conventional $A^{*}$ method. It’s worth noting that while the global planning may exhibit suboptimal outcomes, the local planner effectively mitigates many of its drawbacks, this means that the imperfections are okay and do not need special care when putting the plan into action. In designing the global planner, the paramount considerations include ensuring completeness, accuracy, and reducing computational complexity. Simulations validate that the proposed algorithm in this study satisfactorily fulfills these criteria, making it well-suited for global path planning.
@@ III Experiments
III-A First scenario: The obstacle moves at a constant speed in the vertical (y) direction
@@
III-B Second scenario: The obstacle moves at a constant speed in the vertical (x) direction
@@
III-C Third scenario: The obstacle moves with a constant acceleration
@@
IV Motion planning in the presence of moving obstacles
IV-A First test: Static obstacles
In order to test the performance of the algorithm, the initial values specified in Table II have been used.
TABLE II: Initial conditions for the first test
Description Details
@@ IV-A First test: Static obstacles
$0:[-2,0]$
Initial position of obstacles $1:[2,0]$
$2:[0,0]$
Dynamic conditions of obstacles Completely static
The maximum number of homotopy classes is limited at 5 to control the computational workload. Consequently, 5 trajectories are simultaneously generated, and the one with the lowest cost is chosen. In the fig.5(a) red arrows denote the waypoints for the vehicle to traverse from the origin to the destination. The obstacles are considered as points and the circle around them indicates the safety margin. Obstacles can have any shape and be placed in any position. In order to show how the obstacles affect the path, it has been tried to place the obstacles at a point that has the greatest impact on the path. The final trajectory comprises 81 states, with an estimated time of arrival at 20.14 seconds. The average path computation time is 14.91milliseconds, which is considered efficient for a computer equipped with a dual-core processor running at 1.8 GHz.
Figure 5: (a) Two-dimensional position over time, perspective view, from above, from the front, and from the left in the presence of static obstacles (b) n the presence of constant speed obstacles (c) n the presence of constant acceleration obstacles
IV-B Second scenario: moving obstacles at a constant speed
In order to test the performance of the algorithm, the initial values specified in Table III have been used in this scenario.
TABLE III: Initial conditions for the Second test
Description Details
Initial position of the vehicle $[-4,0]$
Destination location of the vehicle $[4,0]$
$0:[-2,0]$
Initial position of obstacles $1:[2,0]$
$2:[0,0]$
$0:[0.2,0.3]$
Initial speed of obstacles $1:[-0.2,-0.3]$
$2:[0.2,-0.2]$
Dynamic conditions of obstacles Constant velocity
@@ IV-B Second scenario: moving obstacles at a constant speed
IV-C Third test: Accelerated moving obstacles
In order to test the performance of the algorithm, the initial values specified in Table IV have been used.
TABLE IV: Initial conditions for the Third test
Description Details
Initial position of the vehicle $[-4,0]$
Destination location of the vehicle $[4,0]$
$0:[-2,0]$
Initial position of obstacles $1:[2,0]$
$2:[0,0]$
$0:[0.2,0.3]$
Initial speed of obstacles $1:[-0.2,-0.3]$
$2:[0.2,-0.2]$
$0:[-0.01,-0.02]$
Initial speed of obstacles $1:[0.03,0.01]$
$2:[-0.02,0.03]$
Dynamic conditions of obstacles Constant acceleration

### v2

@@ Abstract
Keywords:
Autonomous Vehicles Dynamic obstacles Obstacle avoidance Global planning Local planning Timed elastic band trajectory density.
1 Introduction
@@ 2 Methodology
2.1 Proposed algorithm for global planning
@@ 2.2 $A^{*}$ planner
2.3 Gradient descent method
@@ 2.4 Proposed algorithm for local planning
2.5 Setting the trajectory density
@@
2.6 Obstacle dynamics in local planning
@@
Table 1: System Augmented State
@@
$\hat{x}{\ }_{\text{aug}}(n|n-1)=F_{\text{aug}}(n)\hat{x}{\ }_{\text{aug}}(n-1|n-1)\\ \mathrm{\Sigma}_{\text{aug}}(n|n-1)=F_{\text{aug}}(n)\mathrm{\Sigma}_{\text{aug}}(n-1|n-1)F_{\text{aug}}^{T}(n)\\ +G_{\text{aug}}(n)Q_{\text{aug}}(n)G_{\text{aug}}^{T}(n)\\ y\hat{(n)}=Z_{\text{aug}}(n)-H_{\text{aug}}(n)\hat{x}{\ }_{\text{aug}}(n|n-1)\\ K_{\text{aug}}(n)=[\mathrm{\Sigma}_{\text{aug}}(n|n-1)H_{\text{aug}}^{T}(n)\\ +G_{\text{aug}}(n)T_{\text{aug}}(n)]R_{\text{aug}}^{-1}(n)\\ \hat{x}{\ }_{\text{aug}}(n|n)=\hat{x}{\ }_{\text{aug}}(n|n-1)+K_{\text{aug}}(n)y\hat{(n)}\\ \mathrm{\Sigma}_{\text{aug}}(n|n)=\mathrm{\Sigma}_{\text{aug}}(n|n-1)\\ -K_{\text{aug}}(n)H_{\text{aug}}(n)\mathrm{\Sigma}_{\text{aug}}(n|n-1)$ (10)
@@
The desired dynamic system is modeled using position, speed, and acceleration. Now it is assumed that the acceleration, which is larger than the order, is constant in specific time intervals that are independent of each other and changes at the end of each interval. In other words, the acceleration jumps in each time interval and is modeled as below:
@@
where $\Gamma$ is the system noise gain and $w$ is the desired continuous piece acceleration. The transfer matrix of the system is also defined as follows.
@@
3 Experiments
Fig.2 illustrates the scenario involving a parking lot, showcasing two planning methods: normal $A^{*}$ and gradient descent. Comparing the paths generated by these methods reveals that the gradient descent approach yields a notably smoother and optimal path compared to the conventional $A^{*}$ method. It’s worth noting that while the global planning may exhibit suboptimal outcomes, the local planner effectively mitigates many of its drawbacks, this means that the imperfections are acceptable and do not need special care when putting the plan into action. In designing the global planner, the paramount considerations include ensuring completeness, accuracy, and reducing computational complexity. Simulations validate that the proposed algorithm in this study satisfactorily fulfills these criteria, making it well-suited for global path planning.
@@ 3 Experiments
3.1 First scenario: The obstacle moves at a constant speed in the vertical (y) direction
@@
3.2 Second scenario: The obstacle moves at a constant speed in the vertical (x) direction
@@
3.3 Third scenario: The obstacle moves with a constant acceleration
@@
4 Motion planning in the presence of moving obstacles
4.1 First test: Static obstacles
In order to test the performance of the algorithm, the initial values specified in Table 2 have been used.
The maximum number of homotopy classes is limited at 5 to control the computational workload. Consequently, 5 trajectories are simultaneously generated, and the one with the lowest cost is chosen. In the fig.5(a) red arrows denote the waypoints for the vehicle to traverse from the origin to the destination. The obstacles are considered as points and the circle around them indicates the safety margin. Obstacles can have any shape and be placed in any position. In order to show how the obstacles affect the path, it has been tried to place the obstacles at a point that has the greatest impact on the path. The final trajectory comprises 81 states, with an estimated time of arrival at 20.14 seconds. The average path computation time is 14.91milliseconds, which is considered efficient for a computer equipped with a dual-core processor running at 1.8 GHz. The green paths along by the optimal trajectory are homotopy paths.
Figure 5: (a) Two-dimensional position over time, perspective view, from above, from the front, and from the left in the presence of static obstacles (b) in the presence of constant speed obstacles (c) in the presence of constant acceleration obstacles
Table 2: Initial conditions for all three tests
Description Static Obstacle Obstacle with Constant Speed Obstacle with Constant Acceleration
@@ 4.1 First test: Static obstacles
Initial position of obstacles $0:[-2,0]$ $1:[2,0]$ $2:[0,0]$
Initial speed of obstacles 0:[0,0]:$0$:[0,0]0 : [ 0 , 0 ] $1:[-0.2,-0.3]$ $2:[0.2,-0.2]$
Initial acceleration of obstacles 0:[0,0]:$0$:[0,0]0 : [ 0 , 0 ] $1:[0,0]$ $2:[-0.02,0.03]$
Dynamic conditions of obstacles Completely static Constant velocity Constant acceleration
4.2 Second scenario: moving obstacles at a constant speed
In order to test the performance of the algorithm, the initial values specified in Table 2 have been used in this scenario.
@@ 4.2 Second scenario: moving obstacles at a constant speed
4.3 Third test: Accelerated moving obstacles
In order to test the performance of the algorithm, the initial values specified in Table 2 have been used.